User Tools

Site Tools


We were windstowed at the beginning of this experiment, so were unable to participate in fringe checks. We cabled up what we thought was RCP, but it turned out to be LCP. Therefore useful data started being obtained after 11:11 UT.

We recorded to disks labelled ATNF V006A.

DAS profile - VSOP.PRO

Agilent set to 13.04 GHz, for an effective first LO of 8940 MHz.

2nd LOs set to 637 & 669 MHz.

lbaops/lbanov2008/v252icdlog.txt · Last modified: 2015/12/18 16:38 by